OpenStudy (anonymous):

What is the answer of square root 8x+1=5

OpenStudy (anonymous):

\[\sqrt{8x+1}=5squaring, 8x+1=25\] calculate for x \[also 8 x+1 \ge 0, x \ge - \frac{ 1 }{ 8 }\]